Three children are trying to balance on a seesaw, which consists of a fulcrum rock,

acting as a pivot at the center, and a very light board of 3.6 m long. Two playmates

are already on either end. Boy A has a mass of 50 kg and Girl B and Girl C have

masses of 35 kg and 25 kg respectively.


(a) The Girl C is now sits at a distance x from the fulcrum to balance the seesaw.

Calculate x.


(b) If the weight of the board is to be considered, would distance x increased,

decreased or remain the same as in part (a)?


1
Expert's answer
2021-11-14T17:14:50-0500

"F_1l_1=F_2l_2+F_3l_3\to m_1gl_1=m_2gl_2+m_3gl_3"


"m_1l_1=m_2l_2+m_3l_3\to 50\cdot1.8=35\cdot1.8+25\cdot x \to"


"x=1.08\ (m)" . Answer